OpenStudy (anonymous):

part c) Give the equation for the tangent line in slope-intercept(ax+by=c)form. part a gave me a function r(t)= from 0<=t<=2pi. part b gave me the parametric equation for the tangent line to the curve at t=pi/4. the parametric equation was e^(pi/4)<(sqrt(2)/2),(sqrt(2)/2)+(t)sqrt(2)> i believe that the answer should be x=e^(pi/4)sqrt(2)/2 but i have no idea how to get to it

OpenStudy (turingtest):

did you derive the answer to b yourself? because it looks off to me

OpenStudy (anonymous):

well my professor went over it in class I may have written it down wrong but i dont think so

OpenStudy (anonymous):

i can type out the steps if you would like

OpenStudy (turingtest):

r'(t)<e^t(cost-sint),e^t(sin t+cos t> right?

OpenStudy (anonymous):

correct

OpenStudy (turingtest):

then r'(pi/4)=e^(pi/4)<0,sqrt2>

OpenStudy (anonymous):

right so then the tangent line should be =e^(pi/4)<sqrt2,sqrt2>+te^(pi/4)<0,sqrt2>

OpenStudy (anonymous):

which is e^(pi/4)<(sqrt2)/2,(sqrt2/2)+te^(pi/4)sqrt2>

OpenStudy (turingtest):

ok now I agree, I didn't see which step you were on so write out the formulas for x and y explicitly

OpenStudy (turingtest):

hey you dropped the t in the x component

OpenStudy (anonymous):

i dont think there is one is there?because its te^(pi/4)(0)

OpenStudy (turingtest):

first off r(pi/4)=e^(pi/4)<sqrt2/2,sqrt2/2> you forgot the /2 in one line but you seem to have fixed it in the next

OpenStudy (turingtest):

now the point plus t times the tangent vector is what we need for the equation

OpenStudy (turingtest):

oh shnap you are right, so that's x=sqrt2/2 y=sqrt2/2+e^(pi/4)(sqrt2)t

OpenStudy (turingtest):

x=e^(pi/4)sqrt2/2 y=e^(pi/4)sqrt2/2+e^(pi/4)(sqrt2)t

OpenStudy (anonymous):

ya so i know how to do it when i have 2 t's but now when i have just one

OpenStudy (turingtest):

I don't get it, this should just be a vertical line t does not change x, only y so it should be the line x=e^(pi/4)sqrt2/2 by my reasoning

OpenStudy (anonymous):

ya ok thats what i was told the answer should be but i dont understand how to justify it

OpenStudy (anonymous):

like how do you take the paramertic equations we had and get to that reasoning?

OpenStudy (turingtest):

didn't I? at time t=0 y=x as t changes y does and x does not that must be a vertical line

OpenStudy (turingtest):

you want some more algebra, eh?

OpenStudy (anonymous):

haha idk about that but that does make sense

OpenStudy (anonymous):

thank you

OpenStudy (turingtest):

another way to see it is just by looking at r'(pi/4) it is a vertical direction vector (x-component 0), so again we know it is a vertical line like I said, plugin in t=0 into the parametric equation for the tangent line shows you where y=0, and hence tells you which vertical line it is that is the best I can reason it :P

OpenStudy (turingtest):

welcome!

OpenStudy (turingtest):

ok I thought of a more rigorous way to show it, just find a defined point by plugging in t=0 and you get x=y=e^(pi/4)sqrt2/2 then find the slope by finding\[m={y(b)-y(a)\over x(b)-x(a)}\]for two points t=a and t=b you will of course see that the slope is undefined since x does not depend on t, which means a vertical line passing through the point we found above